OpenStudy (anonymous):

Find the volume of a cone with radius 6 m and height 20 m. Choose one answer. a. 720π m^3 b. 240π m^3 c. 120π m^3 d. 40π m^3

OpenStudy (igreen):

\(\sf V = \dfrac{1}{3} \pi r^2 h\) Plu gin what we know: \(\sf V = \dfrac{1}{3} \pi (6^2)(20)\) Can you simplify that?

OpenStudy (anonymous):

753.6

OpenStudy (igreen):

No, don't simplify pi..we leave that there..just simplify \(\sf\dfrac{1}{3}(6^2)(20)\)

OpenStudy (igreen):

Pi is in the options.

OpenStudy (igreen):

1/3 * 6 * 6 * 20 = ? Simplify that^

OpenStudy (anonymous):

Thanks!